OpenStudy (scorcher219396):

Differentiate (x^2+4x+3)/sqrtx

OpenStudy (phi):

it might be helpful to write \( \sqrt{x} = x^{\frac{1}{2}} \) now you have two choices: 1) use the quotient rule d(u/v) = (u dv - v du)/v^2 or 2) write the x^½ "up top" as x^-½ and use the product rule d(uv) = u dv + v du

OpenStudy (anonymous):

You can also avoid using quotient/product and reduce this to a power rule application: \[\frac{x^2+4x+3}{x^{1/2}}=x^{3/2}+4x^{1/2}+3x^{-1/2}\]